OpenStudy (anonymous):

Simplify 5 log2 k – 8 log2 m + 10 log2 n

OpenStudy (anonymous):

\[5 \log _{2}k-8 \log _{2}m+10\log _{2}n\]

OpenStudy (anonymous):

The 5, k-8, m+10 & n is sepearted not directly together.

myininaya (myininaya):

What does that mean?

myininaya (myininaya):

Is the problem as it appears ?

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[\log_b(a^r)=rlog(a)\] use this rule first

OpenStudy (anonymous):

I don't understand how to.

myininaya (myininaya):

\[5 \log_2k=\log_2(k^5)\] now you try this term: \[8 \log_2m\]

OpenStudy (anonymous):

\[\log _{2}(k ^{m)}\] ?

myininaya (myininaya):

No look at my example and look at the property that I wrote \[rlog_b(a)=\log_b(a^r)\]

myininaya (myininaya):

\[8 \log_2(m)=\log_2(m^8)\]

myininaya (myininaya):